Simplify 7√7 - 2√4 - 2√63
1
Expert's answer
2012-02-07T09:45:32-0500
7 sqrt(7)-2 sqrt(4)-2 sqrt(63)=
7 sqrt(7)-2 sqrt(2*2)-2 sqrt(7*9)=
7 sqrt(7)-2*2-2*3 sqrt(7)=
sqrt(7)-4